Your firm recently paid a dividend of $4 to common stockholders. Dividends are expected to grow at 8% per year for the foreseeable future. The current stock price is $54. New shares could be sold for the same price, but flotation costs would amount to $6 per share.

A $15 million bank line of credit is available with an interest rate of 9%. The firm's tax rate is 34%.

What is the firm’s cost of capital if their capital structure consists of 60% external equity and 40% bank loans?

the answer is 12.58% but need to know how to work this question out
Answer #1

Cost of External Equity = [{D0 * (1 + g)} / (P0 - Flotation Costs)] + g

= [{$4 * (1 + 0.08)} / ($54 - $6)] + 0.08 = 0.09 + 0.08 = 0.17, or 17%

WACC = [wD * kD * (1 - t)] + [wE * kE]

= [0.4 * 9% * (1 - 0.34)] + [0.6 * 17%]

= 2.376% + 10.2% = 12.576%, or 12.58%